# Finding the Fairest Voting System Using Likelihood Analysis

### v2

@@ Abstract
This paper takes a statistical approach to determine which of 10 voting systems, or preference aggregation rules, is the fairest based on their probabilistic likelihood of violating Arrow’s five social choice criteria. The voting systems considered are: Plurality, Borda, Dowdall, Top Two (Plurality Runoff), Instant Runoff, Coombs, Baldwin, Copeland, Pairwise Majority, and Minimax. This paper builds upon the work of Dougherty and Heckelman (2020) by computing violation frequencies for elections with a greater number of alternatives. Elections with up to 50,000 voters and between three and six alternatives are simulated using both Impartial Culture and Impartial Anonymous Culture. The results of these simulations produce new IIA violation likelihoods for each method and show that Pairwise Majority is the most likely to jointly satisfy all five of Arrow’s criteria. Furthermore, of the systems that satisfy transitivity, the Baldwin method is most likely to jointly satisfy all five of Arrow’s criteria in elections with three alternatives. As the number of alternatives increase, the joint satisfaction frequencies decrease rapidly for all systems.
@@ 1 Introduction
Voting systems can be evaluated based on whether they satisfy or violate social choice criteria. This paper takes a statistical approach to determine which voting system is the fairest based on the probabilistic likelihood of violating Arrow’s five social choice criteria.
In total, 10 voting systems are analyzed with respect to five social choice criteria. Specifically, the methods considered are Plurality, Borda, Dowdall, Top Two (Plurality Runoff), Instant Runoff, Coombs, Baldwin, Copeland, Pairwise Majority, and Minimax. Through simulations with up to 50,000 voters and six alternatives, the violation frequencies are computed for each method. To model voters’ preferences, two cultures are used: Impartial Culture (IC) and Impartial Anonymous Culture (IAC). First, voting systems are tested with respect to each individual criteria. Then, for each voting system, Joint-Satisfaction (JS) frequencies, the probability that systems jointly satisfy all five conditions, are computed. These frequencies are used to determine the fairest system.
This paper builds upon the work of previous empirical studies in social choice theory. Compared to prior works, this study computes violation frequencies for elections with a greater number of voters and alternatives than most literature. For the case of three-alternatives elections, the majority of our results align with Dougherty and Heckelman (2020). This paper both extends their analysis for up to six alternatives and finds higher lower bounds for the frequency of IIA violations. The main findings are summarized below:
Result 1: Of the 10 voting systems analyzed, Pairwise Majority is the most likely to jointly satisfy Arrow’s five social choice criteria. In elections with three alternatives and at least 100 voters, Pairwise Majority has Joint-Satisfaction frequencies of over 85%.
Result 2: Of the systems that satisfy transitivity, Baldwin’s Method has the highest Joint-Satisfaction frequencies. The frequencies, however, are less than 20%.
Result 3: As the number of alternatives increases, the Joint-Satisfaction frequencies decrease for each voting method. For all systems except Pairwise Majority, the Joint-Satisfaction frequencies in elections with at least 30 voters and four alternatives are 0%.
The paper proceeds as follows: Sections 2 and 3 detail all required background knowledge. Section 4 contains this study’s contributions and new results. Section 5 describes the experiment set-up and methodology. Section 6 presents the results, and Section 7 discusses the differences in results to past papers. Section 8 concludes the paper and highlights the key points.
@@ 2 Voting Systems
We follow the same assumptions as in Dougherty and Heckelman (2020). We define below the voting systems analyzed in this paper which are a subset of those considered in their paper. In Dougherty and Heckelman (2020), Plurality, Minimax, Top Two, and Instant Runoff are said to be part of Group 1 due to their similar numerical results, and we adopt this convention here.
[…]
@@
Alternatives are ranked based on their first-place vote tally.
@@
Let the election have $n$ alternatives. For each 1st place vote, alternatives receive $n$ points, for each 2nd place vote they receive $n-1$ points, and so on until for an $n$th place vote, alternatives receive one point. To obtain a societal preference order, the alternatives are ranked based on their point total.
@@
Let the election have $n$ alternatives. For each first place vote, alternatives receive 1 point, for a second place vote they receive $\frac{1}{2}$ point, for a third place vote they receive $\frac{1}{3}$ point and so on until for an $n$th place vote, alternatives receive $\frac{1}{n}$ points. The voters are then ranked based on their point tally.
@@
There are two rounds. The first round is a plurality election. If a alternative receives a majority, they win the election. If not, the voters vote again, but can only chose between the two alternatives who receive the most first-place votes in the first round. The alternative with the most first-place votes in the second round wins the election.111 If there is a majority winner in the first round, alternatives are ranked based on their first-place vote tally. If there is no majority winner, the top two alternatives are compared and ranked first and second (or tied first) depending on their relative vote totals in the second round. Then the other alternatives are ranked below them based on their first round vote totals.
@@
In each round, the alternative(s) with the least first place votes is eliminated. The votes of the eliminated alternative(s) are transferred to the alternatives ranked below them. The process repeats until all alternatives are eliminated. Alternatives are ranked based on reverse order of the round they were eliminated.
@@
[…]
@@
This system follows the same procedure as Instant Runoff, but has a different elimination rule. Each round, the alternative with the least points, as calculated by the Borda Count rule, is eliminated.
@@
All alternatives are compared head to head pairwise. For two alternatives $A$ and $B$, if $A$ wins the head to head then $A\succ B$. If $B$ wins the head to head then $B\succ A$. If they tie then $A\sim B$. Based on the comparisons, a transitive ordering is formed if one exists.
@@
All alternatives are compared head to head pairwise. For each comparison, the alternative that wins head to head receives 1 point. If the two alternatives tie they receive $\frac{1}{2}$ point each. After all comparisons are performed, alternatives are ranked based on their point total.
@@
All alternatives are compared head to head. For each alternative, their score is the magnitude of their worst pairwise defeat. Alternatives are ranked in the reverse order of their score (with lower scores giving higher rank).
@@ 2.1 Arrow’s Criteria of Fairness
[…]
@@
  2. 2.
    Unrestricted Domain (UD) - The domain of the voting system can take any preference schedule as input. This means that for every preference schedule, the voting system produces societal preferences between any two alternatives; there are no restrictions on how voters can vote. Note that, for the scope of this paper, this condition does not imply the rankings have to be transitive.
@@
    Transitivity (TR) - The voting method always produces transitive preferences, i.e., the method always yields a societal preference order. Formally, for any three candidates $A$, $B$, and $C$, $A\succeq B$ and $B\succeq C$ implies $A\succeq C$.
@@
    (Pareto) Unanimity (U) - If all voters prefer alternative $A$ over alternative $B$, then in the societal preference order, $A\succ B$.
@@
[…]
@@ 3 Arrow’s Theorem
[…]
Table 1: Criteria satisfied by 10 Voting Systems
Voting System ND UD TR U IIA
@@
Baldwin
[…]
4 New and Expanded Results
The numerical probability of violating each of Arrow’s five conditions is investigated in Dougherty and Heckelman (2020). The study also calculates the likelihood that voting systems violate at least one of Arrow’s conditions. The new contributions of this paper that extend their prior results are:
  1. 1.
    This paper collects more empirical data, analyzing elections for up to 50,000 voters and six alternatives.
  2. 2.
    This paper confirms and collates the results of previous studies, providing extensive data of the frequencies at which voting systems violate criteria.
  3. 3.
    This paper develops a new methodology to test for IIA violations, which finds substantially more violations than previous algorithms, and thus creates a new greatest lower bound for IIA violations.
  4. 4.
    A new metric, Joint-Satisfaction frequencies, is introduced and formalized to test for the fairness of voting systems. This is similar to Dougherty and Heckelman (2020) regarding the likelihood of violating at least one criteria.
The data tables with the results showing percentages of violations for each criterion and Joint-Satisfaction frequencies are located in Appendix A.
@@ 5 Simulations
All computational models were implemented using C++ and Python. The voting systems (Plurality, Borda, Instant Runoff, Coombs, Copeland, Dowdall, Pairwise Majority, Minimax, Top Two, and Baldwin) were tested against Arrow’s criterion (ND, UD, IIA, U, TR) for three to six alternatives.
Ties were handled as follows: For Top Two, there is a possibility of a tie in the first round, where more than two alternatives have equal number of votes. In this case, ties are broken randomly. In all other methods, if alternatives are tied, they have the same rank in the societal preference order.
[…]
For each setting (fixing the criterion, number of voters, number of alternatives, and culture) 10,000 preference schedules were generated. Then the violation frequency, the percentage of times a voting system violates the criterion, was calculated for each of the 10 voting systems. For JS Frequencies, this involved testing voting systems against multiple criteria consecutively. Due to the simulation process, the frequency of violations calculated are numerical probabilities, not theoretical.
@@ 6 Results and Analysis
The following sections details the likelihood of voting systems violating Arrow’s social choice criterion. Note that all voting methods considered satisfy Non-Dictatorship and Unrestricted Domain by definition and as such are excluded from the data.
@@ 6.1 Independence of Irrelevant Alternatives (IIA)
Key Observations in IIA Data
@@
    In elections with at least three alternatives and two voters, Group 1 systems (Plurality, Instant Runoff, Minimax, and Top Two) violate IIA more than 99.99% of the time.
@@
    As the number of voters increases, IIA violations generally increase under IC. For all systems that violate IIA, the frequency of violations approaches 100% as the number of voters increase. For elections with at least 100 voters and three alternatives, IIA is almost never satisfied. However, this trend is not always observed as many voting systems violate IIA at a lower frequency in elections with four voters than elections with three voters. The IAC data does not strictly abide by this trend.
@@
    As the number of alternatives increases, IIA violations increase. In elections with 4 alternatives and at least 100 voters, all systems violate IIA more than 99.7% of the time for both IC and IAC. For elections with at least five alternatives, IIA is almost never satisfied.
Tables 2 and 3 show percentage violations for three-alternative elections under IC and IAC. Figures 1 and 2 plot the IC and IAC data respectively. Note that Pairwise Majority cannot violate IIA, thus the violation frequencies are not computed.
Table 2: Percentage IIA Violations for 3 alternatives (IC Data)
Voting System Number of Voters
3 4 5 10 30 50 100 1000
Plurality 100 100 100 100 100 100 100 100
Instant Runoff 100 100 100 100 100 100 100 100
Top Two 100 100 100 100 100 100 100 100
Minimax 100 100 100 100 100 100 100 100
Copeland 97.35 73.63 99.95 99.99 100 100 100 100
Dowdall 97.35 99.52 92.87 99.79 100 100 100 100
Borda Count 97.35 73.63 92.87 94.79 99.97 99.99 100 100
Coombs 97.35 73.63 92.87 94.79 99.99 100 100 100
Baldwin 97.35 73.63 92.87 94.79 99.96 100 100 100
Figure 1: IIA Violations in elections with three alternatives (IC)
Under IC, in elections with at least 30 voters, the frequency of violating IIA is 100%. In elections with three voters, Dowdall, Copeland, Coombs, Borda, and Baldwin show identical IIA violations.
@@
3 4 5 10 30 50 100 1000
Plurality 100 100 100 100 100 100 100 100
Instant Runoff 100 100 100 100 100 100 100 100
Top Two 100 100 100 100 100 100 100 100
Minimax 100 100 100 100 100 100 100 100
Copeland 89.08 56.66 97.3 97.75 99.8 99.99 100 99.99
Dowdall 89.08 95.03 78.62 91.8 96.72 97.29 96.91 97.03
Borda Count 89.08 56.66 78.62 74.29 84.52 84.54 84.14 83.64
Coombs 89.08 56.66 78.62 74.29 87.4 86.41 85.7 85.52
Baldwin 89.08 56.66 78.62 74.29 83.23 83.38 83.35 82.49
The IAC data has slightly more promising results and there is significant variation between systems. Coombs, Borda, and Baldwin all violate IIA less than 86% of the time, even for a sizable electorate. IIA violations appear to slightly decrease for Baldwin, Borda, and Coombs as the number of voters increases. The percentage violations for 1000 voters is less than that for 100 voters.
Figure 2: IIA Violations in elections with three alternatives (IAC)
Our three-alternative data substantially differs from Dougherty and Heckelman (2020). For Dowdall, Coombs, Copeland, and Baldwin our algorithm reports greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points. These differences are further explained in Section 7.

### v3

# Evaluating Fairness of Voting Systems: Simulating Violations of Arrow’s Conditions

@@ Abstract
This paper builds upon the work of Dougherty and Heckelman (2020) by computing violation frequencies for elections with a greater number of alternatives. These results determine which of the 13 voting systems, or preference aggregation rules, is the fairest based on their probabilistic likelihood of violating Arrow’s social choice criteria. The voting systems considered are: Plurality, Borda, Dowdall, Top Two, Instant Runoff, Coombs, Baldwin, Copeland, Anti-Plurality, Nanson, Ranked Pairs, Pairwise Majority, and Minimax. Elections with up to 10,000 voters and between three and six alternatives are simulated using both Impartial Culture and Impartial Anonymous Culture. The results of these simulations produce new IIA violation frequencies for each method and show that Pairwise Majority is the least likely to jointly violate Arrow’s criteria. As the number of alternatives increases, the joint-violation frequencies increase for each voting method. For all systems except Pairwise Majority, the joint-violation frequencies in elections with at least 30 voters and four alternatives are greater than 98%.
@@ 1 Introduction
Social choice theory studies how individual preferences can be aggregated into collective decisions, often through voting systems. A central challenge in this field is determining whether any such system can satisfy desirable fairness criteria. Kenneth Arrow’s Impossibility Theorem (1963) famously demonstrated that no rank-order voting system can satisfy a set of intuitive fairness conditions - namely, Non-Dictatorship, Unrestricted Domain, Transitivity, Unanimity, and Independence of Irrelevant Alternatives (IIA) - when there are at least three alternatives and two or more voters. While the theorem is negative in nature, it has motivated decades of research focused on evaluating which voting systems are least likely to violate these criteria in practice, especially under probabilistic models of voter behavior (e.g., [22]; [20]).
Building on this literature, this paper uses large-scale simulations to estimate the frequency with which commonly studied voting systems violate Arrow’s criteria. Our work extends the recent empirical analysis of Dougherty and Heckelman (2020), who examined the probability of criterion violations for elections with three alternatives. We extend their results to elections with up to six alternatives, using both the Impartial Culture (IC) and Impartial Anonymous Culture (IAC) probability distributions to generate preference profiles.
Our analysis quantifies the probability that each system satisfies Arrow’s criteria individually and jointly. We introduce a joint-violation metric as a benchmark of fairness: the likelihood that a system violates Arrow’s criteria in a given election. This measure enables meaningful comparison across systems.
The main findings are:
  • Of all systems, Pairwise Majority is the least likely to jointly violate Arrow’s criteria.
  • Among systems that guarantee transitive outcomes, the Ranked Pairs method has the lowest joint-violation frequencies. However, under most settings these frequencies are greater than 70%.
  • As the number of alternatives increases, joint-violation frequencies increases sharply for any number of voters.
This study contributes to the ongoing normative evaluation of voting systems by highlighting how violation frequencies behave in large-scale simulations and across a wider range of alternatives. It provides evidence that while no voting system is perfect, some perform significantly better than others in satisfying foundational fairness criteria. Furthermore, we improve on existing methodology by using a more sensitive algorithm for detecting IIA violations, which yields greater lower bounds and helps reconcile discrepancies in prior results.
The remainder of the paper is structured as follows: Section 2 defines the voting systems under consideration. Section 3 reviews Arrow’s fairness criteria. Section 4 explains the simulation methods. Section 5 presents the main results and Section 6 summarizes our conclusions. Section 7 contains the author’s declaration statements.
@@ 2 Voting Systems
As our work extends the results of Dougherty and Heckelman (2020) to more alternatives, we consider the same set of voting systems defined equivalently below. Note that Plurality, Minimax, Top Two, Instant Runoff, Anti-Plurality, and Nanson are said to be part of Group 1 due to their similar numerical results.
We adopt the following notation: for two alternatives $A$ and $B$, $A\succ B$ indicates society prefers $A$ over $B$, $A\succeq B$ indicates society prefers $A$ at least as much as $B$, and $A\sim B$ indicates society is indifferent between $A$ and $B$.
@@
Alternatives are ranked in descending order based on their first-place vote tally.
Anti-Plurality -
Alternatives receive zero points for every last place vote and 1 point for all other place votes. Alternatives are ranked in descending order based on their point total.
@@
Let the election have $n$ alternatives. For each 1st place vote, alternatives receive $n$ points, for each 2nd place vote they receive n1n-1 points, and so on until for an $n$th place vote, alternatives receive one point. Alternatives are ranked in descending order based on their point total.
@@
Let the election have $n$ alternatives. For each first place vote, alternatives receive 1 point, for a second place vote they receive $\frac{1}{2}$ point, for a third place vote they receive $\frac{1}{3}$ point and so on until for an $n$th place vote, alternatives receive $\frac{1}{n}$ point. Alternatives are ranked in descending order based on their point total.
@@
Begin with plurality. Eliminate all alternatives except the two with the most plurality points. Take a second plurality vote on just those two alternatives and then eliminate the alternative with fewer plurality points. Rank alternatives in reverse order of elimination. (ensuring the alternative(s) eliminated last is most preferred in the social ranking). The remaining alternatives are ranked in descending order.
@@
Begin with plurality. Eliminate the alternative with the fewest plurality points (eliminating all alternatives that tie for fewest points). Repeat the process on the remaining alternatives until all alternatives are eliminated. Rank alternatives in reverse order of elimination (ensuring the alternative(s) eliminated last is most preferred in the social ranking).
@@
Eliminate the alternative(s) ranked last by the most voters (eliminating all alternatives that tie for most last place votes). Repeat the process on the remaining alternatives until all alternatives are eliminated. Rank alternatives in reverse order of elimination (ensuring the alternative(s) eliminated last is most preferred in the social ranking).
@@
Begin with Borda. Eliminate the alternative with the fewest Borda points (eliminating all alternatives that tie for fewest points). Repeat the process on the remaining alternatives until all alternatives are eliminated. Rank alternatives in reverse order of elimination (ensuring the alternative(s) eliminated last is most preferred in the social ranking).
Nanson -
Begin with Borda. Eliminate any alternative(s) at or below the mean of Borda points. Repeat the process on the remaining alternatives until all alternatives are eliminated. Rank alternatives in reverse order of elimination. (ensuring the alternative(s) eliminated last is more preferred in the social ranking).
@@
All alternatives are compared head to head pairwise. For two alternatives $A$ and $B$, if $A$ wins the head to head then $A\succ B$. If $B$ wins the head to head then $B\succ A$. If they tie then $A\sim B$. Rank alternatives based on the pairwise comparisons.
@@
All alternatives are compared head to head pairwise. For each comparison, the alternative that wins head to head receives 1 point. If the two alternatives tie they receive $\frac{1}{2}$ point each. Rank alternatives in descending order of total points.
@@
All alternatives are compared head to head. For each alternative, their score is the magnitude of their worst pairwise defeat. Rank alternatives in reverse order of the magnitude of the worst defeats.
Ranked Pairs -
Begin with pairwise majority and sort pairwise preferences in descending order of vote margins. Lock in each pair starting with the pair that has the greatest vote margin as long as it does not create a cycle.
@@ 2.1 Arrow’s Criteria of Fairness
Following [20, 25, 31], below is a list of Arrow’s voting criteria used in this paper. Let $C=\{1,2,\dots,c\}$ be a finite set of alternatives and $V=\{1,2,\dots,v\}$ be a finite set of voters.
@@
    Unrestricted Domain (UD) - The domain of the voting system is every possible preference schedule.
  2. 2.
@@
    Transitivity (TR) - $\forall a,b,c\in C,a\succeq b,b\succeq c\implies a\succeq c$.
@@
    (Pareto) Unanimity (U) - If every individual prefers $a$ to $b$, then $a\succ b$ in the final societal ranking.
@@
    Independence of Irrelevant Alternatives (IIA) - Society’s preference between any two alternatives only depends on voter’s preferences between these two alternatives [23, 19]. Equivalently, if there are two preference schedules in which each voter has the same relative preference between alternatives $A$ and $B$, the societal preference between $A$ and $B$ must be the same in both preference schedules.
@@ 3 Arrow’s Theorem
Arrow’s theorem shows that no voting system can satisfy all five conditions in a society with at least two voters and at least three alternatives [18]. Table 1 below summarizes which criteria each voting system satisfies assuming an unrestricted domain (UD). A check $\checkmark$ indicates that the voting system always satisfies the criteria while an $\times$ indicates that the system may violate that criterion under an UD. The theoretical properties of voting systems are well studied in literature [23, 28, 30, 27].
Table 1: Criteria satisfied by 13 Voting Systems
Voting System ND TR U IIA
@@
Ranked Pairs
For a formal proof of the above results, see [32].
@@ 4 Simulations
All computational models were implemented using C++ and Python. The voting systems were tested against Arrow’s criterion for three to six alternatives. All data tables can be found in Appendix A.
For Top Two, there is a possibility of a tie in the first round, where more than two alternatives have an equal number of votes. In this case, ties are broken randomly. In all other methods, if alternatives are tied, they have the same rank in the societal preference order.
Two preference distributions, Impartial Culture and Impartial Anonymous Culture [21] are considered. Impartial Culture (IC) assumes each voter is equally likely to select any of the possible preference orders. Impartial Anonymous Culture (IAC) assumes each preference schedule is equally likely to occur. All pseudo-random number generation was done using the Mersenne Twister engine [24] in C++ random library.
For each setting (fixing the criterion, number of voters, number of alternatives, and culture) 10,000 preference schedules were generated. Then the violation frequency, the percentage of times a voting system violates the criterion, was calculated for each of the 13 voting systems. The joint-violation frequencies are computed by calculating the frequency with which systems violate any one of Arrow’s criteria. Due to the simulation process, the frequency of violations calculated are numerical probabilities, not theoretical.
@@ 5 Results and Analysis
The following sections detail the likelihood of voting systems violating Arrow’s social choice criterion. Note that all voting methods considered satisfy Non-Dictatorship and Unrestricted Domain and as such these criteria are excluded from the analysis.
@@ 5.1 Independence of Irrelevant Alternatives (IIA)
The IIA algorithm in Dougherty and Heckelman (2020) works for three-alternative elections. We use a randomized IIA algorithm that enables analysis for elections with any number of alternatives. Our algorithm finds higher number of violations in most cases. The differences in the algorithms are noted in Appendix B and the code is linked in the Declarations section of the paper.
Key Observations
@@
    For a fixed number of voters, as the number of alternatives increases, IIA violations increase. In elections with four alternatives, violation frequencies are greater than 89% under both IC and IAC. For elections with at least five alternatives, IIA is almost always violated.
@@
    For a fixed number of alternatives, as the number of voters increases, IIA violations approach 100% under IC.
@@
    In alignment with Theorem 1 in [20], for at least three alternatives and two voters, Group 1 systems violate IIA 100%.
Tables 2 and 3 show percentage violations for three and four alternative elections under IAC.
@@
3 4 10 30 100 10000
Group 1 100.00 100.00 100.00 100.00 100.00 100.00
Copeland 89.08 56.66 97.75 99.80 100.00 99.98
Dowdall 89.08 95.03 91.8 96.72 96.91 96.93
Borda 89.08 56.66 74.29 84.52 84.14 84.12
Coombs 89.08 56.66 74.29 87.4 85.7 85.91
Baldwin 89.08 56.66 74.29 83.23 83.35 83.14
Ranked Pairs 50.32 56.66 71.64 75.24 75.97 75.87
Table 3: Percentage IIA Violations for 4 Alternatives (IAC Data)
Voting System Number of Voters
3 4 10 30 100 10000
Group 1 100.00 100.00 100.00 100.00 100.00 100.00
Copeland 99.08 99.81 100.00 100.00 100.00 100.00
Dowdall 99.08 99.81 100.00 100.00 100.00 100.00
Borda 99.08 99.81 99.90 99.87 99.93 99.90
Coombs 99.08 99.81 99.91 100.00 100.00 99.99
Baldwin 99.08 99.81 99.76 99.76 99.75 99.74
Ranked Pairs 89.86 89.77 97.23 98.54 98.51 98.66
Under IAC, Ranked Pairs has the lowest violation frequencies. For elections with 10,000 voters and 3 alternatives, Ranked Pairs violates IIA at a frequency of 75.87% for 10000 voters. For the same election, Coombs, Borda, and Baldwin all violate IIA at frequencies between 83% and 86%.
Our three-alternative data substantially differs from Dougherty and Heckelman (2020). For systems not in Group 1, our algorithm reports greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points. These differences are further explained in Appendix B.